What is the solution to the equation ln( x+1) - ln(x) = 2
Answer by stanbon(26291)

What is the solution to the equation ln( x+1) - ln(x) = 2
-----------------------------------
ln[(x+1)/x] = 2
(x+1)x = e^2
x^2 + x - e^2 = 0
-----
x = [-1 +- sqrt(1 -4*-e^2)]/2
-----
x = [ -1 +- sqrt(1+4e^2)]/2
---
x = (-1/2) + (1/2)sqrt(1+4e^2) or x = (-1/2)-(1/2)sqrt(1+4e^2)
==================================================================
Cheers,
Stan H.